# Pitfalls and Outlooks in Using COMET

### v2

@@ 3.5 Score distribution bias [data]
A trivial conclusion is that COMET scores for different translation directions are not comparable. Nevertheless, we caution that the same phenomenon could happen for other features such as the domain, output style, etc. After all, although z-score re-scaling could mitigate this problem, we still encounter two issues. First, it has not been a common practice since WMT22 and it would further contribute to non-objective scores (Knowles, 2021); second, while z-scoring is straightforward for the translation direction, other latent, language-agnostic biases still exist.
@@ 3.6 Domain bias [data]
Zouhar et al. (2024) found that neural metrics like COMET are biased towards particular domains, manifested by worse test performance on unseen domains. Taking inspiration from their work and our discussions on “latent biases” in the previous section, we now raise a question—can we create adversarial hypotheses in test time to exploit the domain bias in training time? Specifically, different domains in the training data can be associated with different score ranges. Hence, by pretending that a test translation is in a particular domain, it might be possible to manipulate its COMET score.
In our experimental setting, to make it explicit to COMET during training, we prepend the target translation with a tag of its domain—in our case, the year the WMT data originated. Note that in each iteration of WMT, systems get higher overall DA scores (e.g. 0.721 for 2019 and 0.749 for 2021). Table 8 illustrates our setup: during training, we tag the scored translation data with its year; during testing, we trial various year prefixes to understand the effect.
@@
One would expect the metric to produce the same score based solely on the translation quality. However, as listed in Table 9, we prove that by merely changing the year tag, we can influence the average score of the test set. During training, the model would be able to observe that 2019 is associated with the worst score and 2021 the best. During test time, the model follows this bias and also extrapolates it to upcoming years where it predicts an improvement in the average DA scores. While the differences might appear small, they are on the same scale as the differences between years in the training data.
@@ 3.7 Lack of multi-reference support [usage]
By design, only one reference can be used as input to the COMET architecture. Whilst one may argue that representing a text reference in a neural space can ease the restriction on word choices, it might still be beneficial to use multiple references to overcome defects in the base embedding model. Therefore, we are curious whether COMET can explicitly and reliably leverage multiple references. We identify three distinct ways in which multiple references have been incorporated in COMET in previous literature (Rei et al., 2020b; Zouhar and Bojar, 2024):
@@
agg: Obtaining an aggregate score per example as follows. A quadruplet of source , hypothesis , reference , and alternative reference is fed to COMET six times in different [src, hyp, ref] arrangements: , as well as . Then, the average score from these six passes is multiplied by where denotes the standard deviation. See Rei et al. (2020b) for further explanation.
@@
Our recommendations for the inclusion of multiple references into COMET or even other neural metrics are aspirational as this topic warrants further investigation. Extending unified pre-training (Wan et al., 2022) with multiple references in the architecture as well as using training objectives more suitable for handling more than one references (Zheng et al., 2018; Fomicheva et al., 2020) can be helpful.
@@ 3.8 Translationese [data]
COMET has been trained with human translations as references and machine translations as hypotheses, where both could be considered “translationese” to a certain extent (Gellerstam, 1986). An intriguing question is whether COMET would be affected by translationese and whether it prefers a translationese to a more natural translation.
| WMT19 EnDe | COMET | Ranking | COMET | Ranking | ||||
| Sys-6823 | 0.845 | 0.840 | 8 | 5 | 0.8207 | 0.8243 | 8 | 4 |
| Sys-6862 | 0.867 | 0.842 | 2 | 2 | 0.8397 | 0.8281 | 2 | 3 |
| Sys-6819 | 0.843 | 0.835 | 9 | 9 | 0.815 | 0.8171 | 9 | 8 |
| Sys-6808 | 0.869 | 0.843 | 1 | 1 | 0.8402 | 0.8284 | 1 | 2 |
| Sys-6974 | 0.866 | 0.842 | 3 | 3 | 0.8384 | 0.8286 | 3 | 1 |
| Sys-6651 | 0.847 | 0.835 | 7 | 8 | 0.8214 | 0.8166 | 7 | 9 |
| Sys-6926 | 0.852 | 0.839 | 4 | 6 | 0.8243 | 0.8218 | 4 | 6 |
| Sys-6763 | 0.851 | 0.841 | 5 | 4 | 0.8229 | 0.8237 | 5 | 5 |
| Online-B | 0.838 | 0.830 | 10 | 10 | 0.8046 | 0.8095 | 10 | 10 |
| Sys-6731 | 0.850 | 0.837 | 6 | 7 | 0.8215 | 0.8204 | 6 | 7 |
@@
We first conduct an experiment to see if the translationese present in the reference would undermine system evaluation with COMET. We consider WMT’s official reference as a standard version and Freitag et al. (2020)’s paraphrased reference as a less translationese reference.444The “paraphrased as-much-as-possible” version. Experiments are carried out under two settings: 1) WMT19 EnDe submissions evaluated by COMET and 2) WMT20 EnDe evaluated by COMET (Rei et al., 2020b). These two settings cover two scenarios—whether the test suite has been used in training the COMET model, or not. A breakdown of COMET scores and rankings for individual systems are listed in Appendix C Tables 13 and 14.
We notice that COMET scores decline dramatically when we switch the reference from the original one to the paraphrased one—aiming to reduce translationese. It means that COMET is indeed sensitive to such changes. Yet interestingly, the overall system ranking in either setting remains rather stable. We find a very high Kendall’s of 0.9827 and 0.9833 on the system rankings in the two settings; pairwise accuracy computed against human judgements also maintained at 0.924. We conclude that translationese in the reference impacts absolute COMET scores but not system ranking. These patterns are consistent regardless of whether the model has been exposed to the test set.
@@
We then attempt to understand if a varying degree of translationese in the system outputs will influence system ranking by COMET. We run COMET and COMET on WMT19 EnDe system outputs as well as their corresponding rephrased outputs against the same source and reference. To acquire paraphrases affordably, we shortlist the top-10 systems’ translations in the previous experiment and we prompt GPT-4o using the prompt outlined in Figure 3.555Accessed gpt-4o-2024-08-06 via API in Aug 2024. We do not feed the source sentence to prevent the model from revising the quality.
We observe that COMET and COMET yield the same system ranking when the original hypotheses are scored. After substituting the hypotheses with their paraphrases, the ranking has changed more under COMET which witnesses much lower pairwise accuracy and Kendall’s compared to COMET. This suggests that COMET is more sensitive to potential changes in the degree of translationese than COMET, but we are unable to measure accuracy or correlation w.r.t. human judgements due to the lack of such.
@@
We note the limitations in the experimental design: 1) we assume that the paraphrased references/hypotheses are as good as the original ones and less translationese, which might be hard to verify; 2) evaluations that anchor to human judgements assume that human evaluators provide assessment solely on the quality (and do not overly insist on adequacy/translationese); 3) our comparison between COMET and COMET only shows that they do not behave the same in dealing with change in translationese without informing us which is more satisfactory. Investigating the above problems would be costly but interesting.
@@ 3.9 Model reporting [usage]
Different COMET models can yield distinct results. Therefore it is important to always specify the specific model for sensible score interpretation and comparison. In this section, we look at to what extent this holds up in scientific literature.
@@
Starting with 1.1K papers that cite one of the COMET papers, we obtain 417 papers from 2021 to 2024 that have an easily accessible PDF version.666Papers in 2020 did not have to report the specific model version as there was only one available at the time.
We parse the PDF text and check if any of the tables contains the string comet.
Within those papers, we check whether the COMET model information is contained in the PDF using a regular expression.777At least one of the following: xcomet\-;
wmt(20|21|22|23)\-comet; wmt\-da\-estimator;
comet[ \-](da|20|21|22|23).
@@
This establishes a lower bound that at least 12% of papers report COMET scores without specific model information.
@@
Always report the specific COMET model, ideally with a link. In addition, it is better to cite the affiliated COMET paper as opposed to the first paper (Rei et al., 2020a), because different checkpoints have variations in training regimes that might be crucial in analysing the evaluation outcome.
@@ 3.10 Discussions on other issues
A statistical hypothesis test merely shows how likely the difference between the average of a model’s scores on a test set and that of another model’s scores on the same test set is caused by random fluctuation. As Kocmi et al. (2024b) examined, the significance between two metric scores can be made arbitrarily high and one can force by using a sizeable test set. However, this tells little about whether this difference is meaningful to a human reader. For this reason, we stress the use of mt-thresholds that converts differences in metric scores to how perceivable they are by human annotators.888kocmitom.github.io/MT-Thresholds
@@
Research nowadays favours experiments on multiple translation directions, as multilingual translation models and large language models become trendy. Recent papers are more often seen to report a (macro-)average COMET score as an aggregate measure to reflect model performance, often in XxEn, EnXx, and AllAll. Whilst indicative to some extent, this custom is not entirely scientific because 1) the score range is inherently distinct for each translation direction and 2) there is no assurance that the scores are on a linear scale. Consequently, an outlying score in a single direction can distort the average, leading to a false claim. Likewise, absolute COMET score differences are not comparable if from different base numbers or directions. We suggest that, when multiple translation directions are of interest, it might be more viable to report the number of wins (against another system) as an aggregation of individual scores.
Optimizing towards COMET.
Owing to its strong correlation with human judgement, there have been recent works that investigate the feasibility of using COMET to further improve translation quality. These strategies either include COMET in a distillation workflow Finkelstein and Freitag (2024); Guttmann et al. (2024), as a data filtering method Peter et al. (2023), as a decoding method Freitag et al. (2022a); Fernandes et al. (2022); Vernikos and Popescu-Belis (2024) or as a training objective Yan et al. (2023).
Nevertheless, as a neural metric, COMET may cease to be a good measure if practitioners over-optimize a system towards it. As Yan et al. (2023) demonstrated, a model trained towards COMET can generate “universal translations” (hallucinations) preferred by COMET regardless of the source sentence. Recently, using COMET-based MBR decoding has become prevalent in shared tasks, e.g. at the IWSLT24 offline translation task by the CMU team. Their system has the highest COMET scores but ranks third in terms of human evaluation in EnDe Ahmad et al. (2024). Similarly, Unbabel-Tower70B at WMT24 also used MBR and dominated all automatic metrics at WMT24 (Kocmi et al., 2024a). We await WMT24’s human evaluation results to understand the implications further.
MBR decoding could be seen as an automatic way to exploit bias in the scoring method, (currently COMET in most cases), so practitioners need to beware of its shortcomings, disclose the use of such, and base system building on multiple (less correlated) metrics (Jon et al., 2023).
@@
Like most MT metrics, COMET is a sentence-oriented metric, but sometimes in the evaluation of MT, we do not provide the systems with sentence-segmented input. This is most often the case in speech translation (ST) where sentence segmentation is treated as part of the task Ahmad et al. (2024). To address the problem of mismatching segmentation between the system output and the reference, a common solution in ST is to re-segment the output using a minimum error rate method (Matusov et al., 2005) in order to force-align it with the reference. Forced alignment can introduce segmentation errors resulting in truncated (thus grammatically incorrect) sentences. There is evidence that COMET, as a metric reliant on sentence embeddings, is more sensitive to these segmentation errors than metrics like BLEU, which rely purely on -grams with no linguistic notion of the sentence (Amrhein and Haddow, 2022). However in their comparison of COMET with human ranking, Sperber et al. (2024) suggested that COMET-based ranking is robust to segmentation errors yet that a “more thorough study of this issue is needed”.
@@ 4 The SacreCOMET Package
To help alleviate problems 3.1 (software version), 3.2 (compute precision), and 3.9 (model reporting), we release a simple package sacrecomet with two functionalities. Given a model name, the first functionality attempts to find the appropriate citation including a link to the paper and a BibTeX: {mdframed} [style=mdframedlisting] ⬇ $ pip install sacrecomet $ sacrecomet cite Unbabel/xcomet-xl https://arxiv.org/abs/2310.10482 @misc{guerreiro2023xcomet, title={xCOMET: Transparent Machine Translation Evaluation through Fine-grained Error Detection}, ... The second functionality semi-automatically detects the local software versions to generate a signature for better reproducibility. Both functionalities can also be run in interactive mode. {mdframed}[style=mdframedlisting]
@@ 5 Conclusion
COMET is currently one of the most powerful automatic metrics or quality estimators for machine translation, consistently achieving the top correlation with human judgement. In comparison to previous metrics, it is a statistical learning model and thus inherits all the related problems in addition to possible technical misconfigurations. We urge practitioners to consider more deeply the use of COMET in non-standard scenarios especially where such training bias might come into play. Beyond these issues, there has been confusion in the literature in reporting the correct COMET model and its correct setting. For improved consistency, we release an easy-to-use tool to assist practitioners.

### v3

@@ 3.5 Score distribution bias [data]
A trivial conclusion is that COMET scores for different translation directions are not comparable. Nevertheless, we caution that the same phenomenon could happen for other features such as the domain, output style, etc. Although z-score re-scaling could mitigate this problem, it has not been a common practice since WMT22 and it would further contribute to non-objective scores (Knowles, 2021). Moreover, while z-scoring is straightforward for the translation direction, other latent, language-agnostic biases still exist.
@@ 3.6 Domain bias [data]
Neural metrics like COMET are biased towards particular domains, manifested by worse test performance on unseen domains (Zouhar et al., 2024a). Taking inspiration from previous work and our discussions on “latent biases”, we now raise a question—can we create adversarial hypotheses at test time to exploit the domain bias in training time? Specifically, different domains in the training data are associated with different score ranges. By pretending that a translation is in a particular domain, it might manipulate its COMET score.
To make it explicit to COMET during training, we prepend the target translation with a tag of its domain—in our case, the year the WMT data originated. Note that in each iteration of WMT, systems get higher overall DA scores (e.g. 0.721 for 2019 and 0.749 for 2021). Table 8 illustrates our setup: during training, we tag the scored translation data with its year; during testing, we trial various year prefixes to understand the effect.
@@
One would expect the metric to produce the same score based solely on the translation quality. However, as shown in Table 9, by merely changing the year tag, we can influence the average score of the test set. During training, the model would be able to observe that 2019 is associated with the worst score and 2021 the best. During test time, the model follows this bias and also extrapolates it to upcoming years where it predicts an improvement in the average DA scores. While the differences appear small, they are on the same scale as the differences between years in the training data.
@@ 3.7 Lack of multi-reference support [usage]
By design, only one reference can be used in COMET. Whilst one may argue that representing a text reference in the neural space can ease the restriction on word choices, it might still be beneficial to use multiple references to overcome defects in the base embedding model. Therefore, we test whether COMET can explicitly and reliably leverage multiple references. We identify three distinct ways in which multiple references have been incorporated in COMET in previous literature (Rei et al., 2020b; Zouhar and Bojar, 2024):
@@
agg: Obtaining an aggregate score per example as follows. A quadruplet of source , hypothesis , reference , and alternative reference is fed to COMET six times in different [src, hyp, ref] arrangements: , as well as . Then, the average score from these six passes is multiplied by where denotes the standard deviation.
@@
Our recommendations for the inclusion of multiple references into COMET or even other neural metrics are aspirational as this topic warrants further investigation. Extending unified pre-training (Wan et al., 2022) with multiple references in the architecture as well as using training objectives more suitable for handling more than one references (Zheng et al., 2018; Fomicheva et al., 2020a) can be helpful.
@@ 3.8 Translationese [data]
COMET has been trained with human translations as references and machine translations as hypotheses, where both could be deemed “translationese” to a certain extent (Gellerstam, 1986).
@@
We first conduct an experiment to see if the translationese present in the reference would undermine system evaluation with COMET. We consider WMT’s official reference as a standard version and Freitag et al. (2020)’s paraphrased reference as a less translationese reference (we use their “paraphrased as-much-as-possible” version). Experiments are carried out under two settings: (1) WMT19 EnDe submissions scored by COMET, and (2) WMT20 EnDe submissions scored by COMET (Rei et al., 2020b). These two settings cover two scenarios—whether the test suite has been used in training the COMET model, or not. A breakdown of COMET scores and rankings for individual systems are listed in Appendix C Tables 12 and 13.
The COMET scores decline dramatically when we switch the reference from the original one to the paraphrased one—aiming to reduce translationese. It means that COMET is indeed sensitive to such changes in the reference. Yet interestingly, the overall system ranking in either setting remains rather stable. We find a very high Kendall’s of 0.9827 and 0.9833 on the system rankings in the two settings; pairwise accuracy computed against human judgements also maintained at 0.924. We conclude that translationese in the reference impacts absolute COMET scores but not system ranking. These patterns are consistent regardless of whether the model has been exposed to the test set.
@@
We then attempt to understand if a varying degree of translationese in the system outputs will influence system ranking by COMET. We run COMET and COMET on WMT19 EnDe system outputs as well as their corresponding rephrased outputs against the same source and reference. To acquire paraphrases affordably, we shortlist the top-10 systems’ translations in the previous experiment and we prompt GPT-4o using the prompt outlined in Figure 3.222We accessed gpt-4o-2024-08-06 via API in Aug 2024. We do not feed the source sentence to prevent the model from revising the quality.
Appendix Table 14 shows that both models yield the same system ranking when the original hypotheses are scored. After substituting the hypotheses with their paraphrases, the ranking has changed more under COMET which witnesses much lower pairwise accuracy and Kendall’s compared to COMET. This suggests that COMET is more sensitive to potential changes in the degree of translationese than COMET.
@@
We note the limitations of our experiments. First, we assume that the paraphrased references are as good as the original ones and less translationese, but we did not verify this when paraphrasing the hypotheses. If the hypothesis quality has been affected, we also assume that the LLM paraphrasing process affects all system outputs in an equal magnitude. Second, the evaluations that anchor to human judgements assume that human evaluators provide assessment solely on the quality and do not overly insist on adequacy/translationese. Third, our comparison between COMET and COMET only shows that they do not behave the same in dealing with change in the degree of translationese in hypotheses.
@@ 3.9 Model reporting [usage]
Different COMET models can yield distinct results. Therefore it is important to always specify the specific model for sensible score interpretation and comparison. In this section, we examine to what extent this holds up in scientific literature.
@@
Starting with 1100 papers that cite one of the COMET papers, we obtain 417 papers from 2021 to 2024 that have an easily accessible PDF version.333Papers in 2020 did not have to report the specific model as there was only one available at the time. Further, we acknowledge potential bias to only papers with available PDFs.
We check if any of the tables contains the string comet.
Within those papers, we check whether the COMET model information is contained in the PDF using a regular expression.444Case-insensitive: “comet[ \-](da|20|21|22|23)|wmt
(20|21|22|23)\-comet|xcomet\-|wmt\-da\-estimator”
@@
This establishes that at least 12% of papers report COMET scores without specific model information.
@@
Always report the COMET version, ideally with a link. Also, cite the affiliated COMET paper as opposed to the first paper (Rei et al., 2020a), because different checkpoints have variations in training regimes that might be crucial in analysing the evaluation outcome.
@@ 3.10 Discussions on other issues
Statistical hypothesis test merely shows how likely the difference between the average of two model’s scores on the same test set is caused by random fluctuation. Kocmi et al. (2024c) shows the significance of a difference between two metric scores can be made arbitrarily high and one can force by using a sizeable test set. This tells little about whether this difference is meaningful to a human reader. For this reason, we stress the use of mt-thresholds that converts differences in metric scores to how perceivable they are by human annotators.555kocmitom.github.io/MT-Thresholds
@@
Research nowadays favours experiments on multiple translation directions, as multilingual translation models and large language models become trendy. Recent papers are more often seen to report a (macro-)average COMET score as an aggregate measure across many directions, usually in XxEn, EnXx, and AllAll. Whilst indicative, this is not entirely scientific because (1) the score range is inherently distinct for each translation direction and (2) there is no assurance that the scores are on a linear scale. Consequently, an outlying score in a single direction can distort the average, leading to a false claim. Likewise, absolute COMET score differences are not comparable if from different base numbers or directions. We suggest that, when multiple translation directions are of interest, in addition to averaged scores, practitioners can report the number of wins (against another system) as another aggregation of individual scores.
Optimizing to COMET.
Owing to COMET’s strong correlation with human judgement, recent works investigate the feasibility of using it in translation modelling directly. These strategies either include COMET in a distillation workflow Finkelstein and Freitag (2024); Guttmann et al. (2024), as a data filtering method Peter et al. (2023), as a decoding method Freitag et al. (2022a); Fernandes et al. (2022); Vernikos and Popescu-Belis (2024) or as a training objective Yan et al. (2023).
Nevertheless, COMET may cease to be a good measure if practitioners over-optimize a system towards it. As Yan et al. (2023) demonstrated, a model trained towards COMET can generate “universal translations” (hallucinations) preferred by COMET regardless of the source sentence. Recently, using COMET-based MBR decoding has become prevalent in shared tasks. For example, Unbabel-Tower70B at WMT24 also used MBR and dominated all automatic metrics but not so much under human evaluation (Kocmi et al., 2024a, b). MBR decoding could be seen as an automatic way to exploit bias in the scoring method, (currently COMET in most cases), so practitioners need to be aware of its shortcomings, disclose the use of such, and base system building on multiple (less correlated) metrics (Jon et al., 2023).
A novel issue is using automated metrics in human evaluation (Zouhar et al., 2024b) that collects data for metric training. This might create a similar effect as translationese in machine translation. The data could be biased by the particular quality estimator that is assisting annotators in the data collection process.
@@
Like most MT metrics, COMET works at the sentence level, but sometimes sentence-segmented input is not available. This is often the case in speech translation (ST) where sentence segmentation is treated as part of the task Ahmad et al. (2024). To address the problem of mismatching segmentation between the system output and the reference, a common solution in ST is to re-segment the output using a minimum error rate method (Matusov et al., 2005) in order to force-align it with the reference. Forced alignment can introduce segmentation errors resulting in truncated (thus grammatically incorrect) sentences. There is evidence that COMET, as a metric reliant on sentence embeddings, is more sensitive to segmentation errors than string-based metrics, like BLEU, which rely purely on -gram overlaps with no linguistic notion of a sentence (Amrhein and Haddow, 2022). In a recent comparison of COMET with human ranking, Sperber et al. (2024) suggested that COMET-based ranking is robust to segmentation errors but that a “more thorough study of this issue is needed”.
Other metrics.
Our work focused on COMET, the current most popular family of MT metrics. Nonetheless, our recommendations could apply to other neural metrics, like MetricX-23 (Juraska et al., 2023), because many issues we outlined are due to their statistical learning nature. Even beyond this, metric reporting and score interpretation in practice, e.g. software usage or averaging across multiple directions, can be problematic for string-matching metrics like BLEU or chrF too.
@@ 4 The SacreCOMET Package
To help alleviate problems in Sections 3.1 (software version), 3.2 (compute precision), and 3.9 (model reporting), we release a simple package sacreCOMET with two functionalities. Given a model name, the first functionality attempts to find the appropriate citation including a link to the paper and a BibTeX:
[style=mdframedlisting]
The second functionality semi-automatically detects the local software versions to generate a signature for better reproducibility. Both functionalities can also be run in interactive mode.
[style=mdframedlisting]
5 Future Work on Learned Metrics
Fixing data bias: Learned metrics are sensitive to the training data distribution. Future metrics should aim to reduce the bias caused by the data selection process such that they are applicable to a range of MT systems.
Interpretability across languages: Currently, practitioners cannot compare, or pedantically, aggregate scores in different translation directions. It would be useful to unify the scores to a single scale that can be interpreted independent of the language (similar to Kocmi et al., 2024c), e.g. to indicate X% of segments are production-ready.
Inference speed: Learned metrics are getting better but at the cost of bulky models and increased inference time. These overheads should be taken into account when developing new models, such as the work of Rei et al. (2022b).
Representations: Current COMET models are built upon off-the-shelf multilingual encoder models which are likely trained on human-written texts. However, this could bring in a domain mismatch—when translation hypotheses act as the input to metric models, they are not human-written but machine-translated.
Robustness: Metrics should have the correct behaviour even in corner cases, be it empty output or incorrect language. Mapping all inputs (Amrhein et al., 2022, inter alia), including partial or adversarial ones, evaluating the metrics, and coming up with methods to make them more robust would increase the metrics’ adoption and trust.
Built-in QE: In production, machine translation and quality estimation are commonly two different processes. In many applications, however, a single QE model is used for a single MT model. Quantifying how much is QE adaptation to a particular MT model useful is beneficial for a holistic understanding of QE metrics. Further, proposing methods for supervised quality estimation built into the MT could ease industry adoption. Beyond the work of Tomani et al. (2024), this remains largely unexplored.
Noise-aware training: Human annotations are notoriously noisy. At the scale of WMT data, poor-quality annotations are unavoidable. The inter-annotator agreement for even robust annotations, such as ESA, remains low at . The effect of data quality on learned metrics is so far unknown and methods for noise/uncertainty-aware training are under-studied.
@@ 6 Conclusion
COMET is currently one of the most powerful automatic metrics/quality estimators for machine translation, consistently achieving the top correlation with human judgement. In comparison to previous metrics, it is a statistical learning model and thus inherits all the related problems in addition to possible technical misconfigurations. We urge practitioners to consider more deeply the use of COMET in non-standard scenarios especially where such training bias might come into play. Beyond these issues, there has been confusion in the literature in reporting the correct COMET model and its correct setting. For improved consistency, we release an easy-to-use tool to assist practitioners.